# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 1 Introduction
This motivates our present study. As to the rare event simulation method, we use a method alternate to action minimisation termed Adaptive Multilevel Splitting (AMS) [Cérou and Guyader, 2007, Cérou et al., 2011], to compute the rare transitions from laminar baseflow to turbulence in plane Couette flow, in domains of increasing sizes at transitional Reynolds numbers. By decreasing the variance of the forcing, proportional to the energy injection rate, we will investigate the possible concentration of reactive trajectories and whether this trajectories are structured by an instanton. Using several noise spectra, we will test whether forcing directly the most responsive structure is a sine qua non condition to travel to turbulence, while increasing the system size will highlight possible spatial localisation of the growing turbulence. We thus present this study in the following manner. We first present the forced Navier–Stokes equations (§ 2.1), then the version of Adaptive Multilevel Splitting that is used to compute the trajectories (§ 2.2), including the details of the set up of AMS computations and additional diagnostics used to study the trajectories such as bisections. We describe the computed paths in section 3, first visually in relevant examples (§ 3.1), then statistically, on the whole collection of computed path (§ 3.2). We finally study the effect of the energy injection rate, on the transition rate (§ 4.1) and on the trajectories duration (§ 4.2). We will discuss these results in view of the literature in the conclusion (§ 5).
@@ 2.1 Forced Navier–Stokes equations
Figure 1: A picture of transition to turbulence in plane Couette flow indicating the type of flow obtained as the Reynolds number is increased. Plane Couette flow is sketched on the left: we indicate the two walls moving at velocities ±$\pm U$, the domain size $L_{x}$, $L_{z}$ and $h$ as well as the coordinate system. For Reynolds numbers larger than a $R_{\rm g}\simeq 325$, laminar-turbulent coexistence can be sustained, as indicated by the colour levels of kinetic energy $\frac{1}{2}\mathbf{u}^{2}$ in the horizontal midplane $y=0$ obtained by means of direct numerical simulations in a domain of size $L_{x}\times L_{z}=110\times 32$ at $R=370$ (centre). The coloured frames indicate the domain sizes considered here (cyan: $L_{x}\times L_{z}=6\times 4$, blue $L_{x}\times L_{z}=18\times 12$, green $L_{x}\times L_{z}=36\times 24$). As the Reynolds number is further increased, wall turbulence can invade the whole domain, as illustrated by the colour levels of kinetic energy $\frac{1}{2}\mathbf{u}^{2}$ in the horizontal midplane $y=0$ obtained by means of direct numerical simulations in a domain of size $L_{x}\times L_{z}=110\times 32$ at $R=370$ (right). The visualisations are performed using numerical results originally performed for [Rolland, 2015], and the whole figure was originally created for [Rolland, 2022].
@@ 2.1.1 Governing equations
We consider plane Couette flow, the flow between two parallel walls separated by a distance $2h$ moving at velocities ±$\pm U$ (sketched in figure 1, left). We use our standard notations [Rolland, 2022] so that $\mathbf{e_{x}}$ is the streamwise direction, $\mathbf{e_{y}}$ is the wall normal direction and $\mathbf{e_{z}}$ is the spanwise direction. Velocities are made dimensionless with $U$, lengths are made dimensionless with $h$ and times are made dimensionless with $h/U$. The main dimensionless control parameter is the Reynolds number $R=\frac{hU}{\nu}$ where $\nu$ is the kinematic viscosity. The flow is also controlled by the dimensionless domain sizes $L_{x}$ and $L_{z}$.
@@
Using tensorial notations, we denote each component of $\mathbf{u}$ by $u_{l}$, where subscript $l$ and $j$ stands for the component $\textbf{e}_{x}$, $\textbf{e}_{y}$ and $\textbf{e}_{z}$ (respectively $l,j=1,2,3$). The forcing $\mathbf{f}$ has zero ensemble average and has in plane spatial correlations. we will see that its energy injection rate is controlled by $\beta$. Before giving further details on $\mathbf{f}$, we need to present the numerical discretisation that we use. The Navier–Stokes equations are discretised in space using $N_{x}$ and $N_{z}$ dealiased Fourier modes in the $x$ and $z$ direction (so that $\frac{3}{2}N_{x}$ and $\frac{3}{2}N_{z}$ modes are used in total) and $N_{y}$ Chebyshev Polynomials in the $y$ direction. This discretisation and the time integration are performed using a code based on channelflow, by J. Gibson [Gibson et al., 2008]. We will use a fixed time step $\Delta t$ for more control on our numerical procedure. For each system size, Reynolds number and resolution, the value of $\Delta t$ is chosen so that the CFL criterion is respected. We will present results of simulations performed in domains of size $L_{x}\times L_{z}=36\times 24$, where we used $\frac{3}{2}Nx=192$, $\frac{3}{2}N_{z}=144$, $N_{y}=27$ and $\Delta t=0.05$, $L_{x}\times L_{z}=18\times 12$, where we used $\frac{3}{2}Nx=96$, $\frac{3}{2}N_{z}=72$, $N_{y}=27$ and $\Delta t=0.05$. Finally, in MFU type domains of size $L_{x}\times L_{z}=6\times 4$, we considered four resolutions:
  • $\frac{3}{2}N_{x}=32$, $\frac{3}{2}N_{z}=24$, $N_{y}=27$ and $\Delta t=0.05$, that we termed standard resolution,
  • $\frac{3}{2}N_{x}=64$, $\frac{3}{2}N_{z}=48$, $N_{y}=47$, and $\Delta t=0.0125$, that we termed high resolution,
  • $\frac{3}{2}N_{x}=64$, $\frac{3}{2}N_{z}=48$, $N_{y}=27$ and $\Delta t=0.05$,
  • $\frac{3}{2}N_{x}=128$, $\frac{3}{2}N_{z}=96$, $N_{y}=35$ and $\Delta t=0.025$,
to test numerical convergence with spatial resolution of the simulations. We will comment on the effect of resolution throughout the text. We present the domain sizes in figure 1, centre, using concentric frames.
@@
$\langle f_{l}(\textbf{x},t)f_{l}(\textbf{x}^{\prime},t^{\prime})\rangle=\delta(t-t^{\prime})C_{l}(x-x^{\prime},z-z^{\prime})\delta(y-y^{\prime})\,,\,\hat{C}_{l}(n_{x},n_{z})=\Gamma_{l,n_{x},n_{z}}\,.$ (3)
@@
The force $\mathbf{f}$ is white (as indicated by the delta correlations) in time and in the wall normal direction. Our numerical simulations use a finite number $N_{y}$ of Chebyshev modes in the wall normal direction. This means that with a red spectrum in $x$ and $z$, the force $\mathbf{f}$ injects a finite amount of energy. The correlation function $C_{l}(x-x^{\prime},z-z^{\prime})$ in the streamwise and spanwise direction on component $l$ is defined using its Fourier transform. Said transform is expressed using a shape factor $\Gamma_{l,n_{x},n_{z}}$, where $n_{x}$ stands for the streamwise wavenumber, $n_{z}$ stands for the spanwise wavenumber. We have followed two different options as to the details of the forcing, presented in details in appendix A. One type of noise is forcing selected components (either a single components or two or all three components). In that case we have that different components are decorrelated $\langle f_{l}f_{j}\rangle=0$ if $l\neq j$. Another type of forcing is divergence free and forces components $\mathbf{e}_{x}$ and $\mathbf{e}_{z}$: it is the curl of a vector potential along $\mathbf{e}_{y}$. In that case components $\mathbf{e}_{x}$ and $\mathbf{e}_{z}$ are forced and a correlation can exists between these two components of the forcing. Numerically, we chose to add the forcing after the predictor-corrector stage of time integration, in line with the formulation of an Itô Process. We checked a posteriori that incompressibility was enforced up to the same numerical precision as in unforced simulations. Note that in studies of action minimisation in plane Poiseuille flow, both an unconstrained noise white in time and space [Wan, 2013] and a noise defined as a Wiener process on the space of divergent free vector fields [Wan et al., 2015, Wan and Yu, 2017], still white in time, were considered. Linear [Farrell and Ioannou, 1993b] and weakly non linear [Ducimetière et al., 2022] studies of flows used noise with little constraint on spectrum shape and component correlations. In the large majority of cases, we will not directly force the spatial scales expected in the flow of order $\mathcal{O}(1)$ or larger: we force below a corresponding cut-off scale $L_{c}$. The value of this scale will be specified in each forcing case. What we do is that we let the flow non linearly select its scales from energy injected solely at moderately small scales (some comparison with forcing all large scales have been performed). Both types of forcing typically give a flat spectrum for the forcing at intermediate wave numbers and after a wavenumber cut-off $N_{r}$, we will use a decaying spectrum for the forcing at larger wavenumbers. We display an example of shape factor $\Gamma_{x,n_{x},n_{z}}$ in the divergence free case in figure 2 (a) and $\Gamma_{z,n_{x},n_{z}}$ in figure 2 (b), we we chose $L_{c}=1.0$. As a comparison, the spectra of $u_{x}$ and $u_{z}$ computed in the midplane $y=0$ for transitional turbulence in a MFU type domain at the largest resolution are displayed in figure 2 (c) and figure 2 (d) respectively. We can notice that these two components have their intense modes in the wavenumber range $\left|\frac{2\pi n_{x}}{L_{x}}\right|\leq 3$, $\left|\frac{2\pi n_{z}}{L_{z}}\right|\leq 6$, where the forcing is absent. The forcing starts to act in a wavenumber range where the natural Fourier modes of the velocity field are at least one order of magnitude smaller than their maximum.
@@ 2.1.2 Diagnostics of the flow
$E_{k,l}=\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{u_{l}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z\,,\,E_{k,y-z}=\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{u_{y}^{2}+u_{z}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z\,.$ (5)
The kinetic energy in the streamwise component $E_{k,x}$ will be the most used. The energy $E_{k,x}$ roughly quantifies the amount of energy contained in the velocity streaks, while the energy $E_{k,y-z}$ roughly quantifies the amount of energy contained in the streamwise vortices, which are the two main coherent structures in transitional wall flows. We finally used a turbulent fraction $F$ to observe the possible spatial localisation of turbulence. This indicator is computed in the following manner: we spatially average the kinetic energy contained in components $u_{y}$ and $u_{z}$ in $\tilde{N_{z}}$ streamwise bands of width $\delta z=L_{z}/\tilde{N_{z}}=2.0$, to define
@@
We now introduce a parameter alternative to $\beta$ which is possibly better suited to follow the transitions: the energy injection rate. For this purpose, we multiply equation (2) by $u_{l}$, sum over the three components and simplify using incompressibility. We use Itô’s Lemma to manage this non linear change of variables. We perform a spatial average and an ensemble average with respect to the forcing noise $\mathbf{f}$. This yields the average energy budget
@@
in the simulation. In this energy budget (Eq. (7)), the extraction is performed by lift-up, in the streamwise direction, thus involving $u_{x}$ and $u_{y}$ and the wall normal gradient of the laminar baseflow equal to $1$. Dissipation is performed by viscosity. In the case of an unforced turbulent plane Couette flow, these two terms compensate one another after time average. For small energy injection rate $\epsilon$, it is generally expected that the flow remains close to laminar and that equation (2) can be treated linearly [Farrell and Ioannou, 1993b] or could be processed using a weakly non linear approach [Ducimetière et al., 2022]. However, this behaviour is conditioned on no transition happening and is thus not really represented by the actual ensemble average in equation (7). We will see in the rest of this text that even in that case, transitions can occur. This means that starting from the initial condition $\langle E_{k}\rangle(t=0)=0$ even with a small $\epsilon$, the solution of equation (7) would display a very slow but steady increase of $\langle E_{k}\rangle(t>0)$ until the flow becomes turbulent. Let us discuss how transition would be represented in this energy budget. We would expect that $\epsilon$ would act against dissipation and could maintain some energy extraction (that would not be sustainable without energy injection, regardless of the spectrum shape). It is possible that a large part of the energy transfers would not be visible in this equation as they would be erased by the spatial averages. We would also expect that the larger $\epsilon$ is, the more efficiently transition is forced, similarly to the dependence on amplitude of perturbations to boundary layers [Spangler and Wells Jr, 1968, Rigas et al., 2021]. The energy injection rate is controlled by two parameters whose effect on the build-up can be presented independently. Indeed, from equation (8), we deduce that $\beta$ and the spectrum shape are the two independent control parameters of this energy injection rate $\epsilon$. Thus, on the one hand, decreasing $\beta$ increases $\epsilon$, and on the other hand, increasing the sum of the $\Gamma_{l,n_{x},n_{z}}$ also increases $\epsilon$: in both cases, we force the flow more intensely. In our definitions (§ A), these shape factors $\Gamma$ will be normalised and will decrease at large wavenumbers. Thus the sum will be larger if that decrease starts at larger wave numbers and if that decrease is slow. Note that we cannot rule out that there could be effect of the precise shape of the spectrum. These complex effects would be hidden in the way energy is transferred from the forced scales to the most receptive scales. There can also be effects of a direct forcing of scales versus energy transfer. As a final remark, equation (8) reminds that in order to have a finite or even vanishing energy injection rate in cases where the flow is forced with a noise white in space (see [Wan, 2013]), one would need to take the limit of $\beta\rightarrow 0$ first and $N_{x},N_{y},N_{z}\rightarrow\infty$ next (or use specific relation between those parameters). With our spectrum shape, we keep the energy injection rate finite, similarly to what is done in study of linear response [Farrell and Ioannou, 1993b] and non linear response [Ducimetière et al., 2022] to stochastic forcing. This physical argument leads to the same mathematical condition for well posedness of the Freidlin–Wentzell Large Deviation principle in two dimensional Navier–Stokes equations as noted by [Wan et al., 2015] and demonstrated by [Brzeźniak et al., 2015].
@@
Note that a similar energy budget is also commonly used in the study of transition to turbulence [Kawahara and Kida, 2001, Gibson et al., 2008, Kreilos and Eckhardt, 2012]. This time, it is obtained for the total spatially averaged kinetic energy $E_{\rm tot}$. For this matter, we apply Itô’s Lemma to the Navier–Stokes equations for the total flow $\mathbf{u}+y\mathbf{e}_{x}$, followed by spatial and ensemble average gives us
$\frac{dE_{\rm tot}}{dt}=\frac{1}{R}(I-D)+\epsilon\,,\,E_{\rm tot}=\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{(u_{x}+y)^{2}+u_{y}^{2}+u_{z}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z$
$I=1+\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{z=0}^{L_{z}}\left(\frac{\partial u_{x}}{\partial y}(y=+1)+\frac{\partial u_{x}}{\partial y}(y=-1)\right)\,{\rm d}x{\rm d}z\,,$
@@
where $I$ is the total energy injection at the wall and $D$ is the total dissipation, written using total vorticity because of incompressibility, periodic and no slip boundary conditions. As such, both are rescaled by their value in the laminar flow where $\mathbf{u}=\mathbf{0}$ and thus $I=D=1$. This pair of variable is convenient to observe transition in wall flows. The line $I=D$ has a special importance because all invariant states lie on it (since $\frac{dE_{\rm tot}}{dt}=0$ on them), while limit cycles lie over it (since the time average over a period of $\frac{dE_{\rm tot}}{dt}$ is 00 on them). Meanwhile turbulent flow is close to $I=D=3$ [Kawahara and Kida, 2001, Gibson et al., 2008].

### v2

@@ 1.3 Outline of the article
This motivates our present study. On the one hand, we would like to propose a method to study transition that can emulate experiments Spangler and Wells Jr, (1968); Liu et al., (2021). On the other hand, we would like to study multistability forced by noise in the case of a very complex separatrix. As to the rare event simulation method, we use a method alternate to action minimisation termed Adaptive Multilevel Splitting (AMS) Cérou and Guyader, (2007); Cérou et al., (2011), to compute the rare transitions from laminar baseflow to turbulence in plane Couette flow, in domains of increasing sizes at transitional Reynolds numbers. By decreasing the variance of the forcing, proportional to the energy injection rate, we will investigate the possible concentration of reactive trajectories and whether this trajectories are structured by an instanton. Using several noise spectra, we will test whether forcing directly the most responsive structure is a sine qua non condition to travel to turbulence, while increasing the system size will highlight possible spatial localisation of the growing turbulence. We thus present this study in the following manner. We first present the forced Navier–Stokes equations (§ 2.1), then the version of Adaptive Multilevel Splitting that is used to compute the trajectories (§ 2.2), including the details of the set up of AMS computations and additional diagnostics used to study the trajectories such as bisections. We describe the computed paths in section 3, first visually in relevant examples (§ 3.1), then statistically, on the whole collection of computed path (§ 3.2). We finally study the effect of the energy injection rate, on the transition rate (§ 4.1) and on the trajectories duration (§ 4.2). We will discuss these results in view of the literature in the conclusion (§ 5).
@@ 2.1 Forced Navier–Stokes equations
Figure 1: A picture of transition to turbulence in plane Couette flow indicating the type of flow obtained as the Reynolds number is increased. Plane Couette flow is sketched on the left: we indicate the two walls moving at velocities ±$\pm U$, the domain size $L_{x}$, $L_{z}$ and $h$ as well as the coordinate system. For Reynolds numbers larger than a $R_{\rm g}\simeq 325$, laminar-turbulent coexistence can be sustained, as indicated by the colour levels of kinetic energy $\frac{1}{2}\mathbf{u}^{2}$ in the horizontal midplane $y=0$ obtained by means of direct numerical simulations in a domain of size $L_{x}\times L_{z}=110\times 32$ at $R=370$ (centre). The coloured frames indicate the domain sizes considered here (cyan: $L_{x}\times L_{z}=6\times 4$, blue $L_{x}\times L_{z}=18\times 12$, green $L_{x}\times L_{z}=36\times 24$). As the Reynolds number is further increased, wall turbulence can invade the whole domain, as illustrated by the colour levels of kinetic energy $\frac{1}{2}\mathbf{u}^{2}$ in the horizontal midplane $y=0$ obtained by means of direct numerical simulations in a domain of size $L_{x}\times L_{z}=110\times 32$ at $R=370$ (right). The visualisations are performed using numerical results originally performed for Rolland, (2015), and the whole figure was originally created for Rolland, (2022).
@@ 2.1.1 Governing equations
We consider plane Couette flow, the flow between two parallel walls separated by a distance $2h$ moving at velocities ±$\pm U$ (sketched in figure 1, left). We use our standard notations Rolland, (2022) so that $\mathbf{e_{x}}$ is the streamwise direction, $\mathbf{e_{y}}$ is the wall normal direction and $\mathbf{e_{z}}$ is the spanwise direction. Velocities are made dimensionless with $U$, lengths are made dimensionless with $h$ and times are made dimensionless with $h/U$. The main dimensionless control parameter is the Reynolds number $R=\frac{hU}{\nu}$ where $\nu$ is the kinematic viscosity. The flow is also controlled by the dimensionless domain sizes $L_{x}$ and $L_{z}$.
@@
Using tensorial notations, we denote each component of $\mathbf{u}$ by $u_{l}$, where subscript $l$ and $j$ stands for the component $\textbf{e}_{x}$, $\textbf{e}_{y}$ and $\textbf{e}_{z}$ (respectively $l,j=1,2,3$). The forcing $\mathbf{f}$ has zero ensemble average and has in plane spatial correlations. We will see that its energy injection rate is controlled by $\beta$. Before giving further details on $\mathbf{f}$, we need to present the numerical discretisation that we use. The Navier–Stokes equations are discretised in space using $N_{x}$ and $N_{z}$ dealiased Fourier modes in the $x$ and $z$ direction (so that $\frac{3}{2}N_{x}$ and $\frac{3}{2}N_{z}$ modes are used in total) and $N_{y}$ Chebyshev Polynomials in the $y$ direction. This discretisation and the time integration are performed using a code based on channelflow, by J. Gibson Gibson et al., (2008). We will use a fixed time step $\Delta t$ for more control on our numerical procedure. For each system size, Reynolds number and resolution, the value of $\Delta t$ is chosen so that the CFL criterion is respected. We give the numerical resolutions used for our simulations in the three considered domain sizes $L_{x}\times L_{z}=6\times 4$, $18\times 12$ and $36\times 24$ in table 1. Note that we have considered several resolutions for the smallest domain $L_{x}\times L_{z}=6\times 4$ to test numerical convergence with spatial resolution of the simulations. We will comment on the effect of resolution throughout the text. We present the domain sizes in figure 1, centre, using concentric frames.
$L_{x}\times L_{z}$ $\frac{3}{2}N_{x}$ $\frac{3}{2}N_{z}$ $N_{y}$ $\Delta t$ resolution type $36\times 24$ 192 144 27 0.05 $18\times 12$ 96 72 27 0.05 $6\times 4$ 32 24 27 0.05 standard resolution $6\times 4$ 64 48 47 0.0125 high resolution $6\times 4$ 64 48 27 0.05 $6\times 4$ 128 96 35 0.025
Table 1: Table listing the domain sizes $L_{x}\times L_{z}$ considered in simulations, along with the number of dealiased Fourier modes in the streamwise and spanwise directions $\frac{3}{2}N_{x}$, $\frac{3}{2}N_{z}$, the number of Chebyshev modes $N_{y}$ and the time step $\Delta t$. Two specific resolutions are distinguished in the smaller domain of size $L_{x}\times L_{z}=6\times 4$.
@@
$\langle f_{l}(\textbf{x},t)f_{l}(\textbf{x}^{\prime},t^{\prime})\rangle=\delta(t-t^{\prime})C_{l}(x-x^{\prime},z-z^{\prime})\delta(y-y^{\prime})\,,\\ ,\hat{C}_{l}(n_{x},n_{z})=\Gamma_{l,n_{x},n_{z}}\,.$ (3)
@@
The force $\mathbf{f}$ is white (as indicated by the delta correlations) in time and in the wall normal direction. Our numerical simulations use a finite number $N_{y}$ of Chebyshev modes in the wall normal direction. This means that with a red spectrum in $x$ and $z$, the force $\mathbf{f}$ injects a finite amount of energy. The correlation function $C_{l}(x-x^{\prime},z-z^{\prime})$ in the streamwise and spanwise direction on component $l$ is defined using its Fourier transform. Said transform is expressed using a shape factor $\Gamma_{l,n_{x},n_{z}}$, where $n_{x}$ stands for the streamwise wavenumber, $n_{z}$ stands for the spanwise wavenumber. We have followed two different options as to the details of the forcing, presented in details in appendix A. One type of noise is forcing selected components (either a single components or two or all three components). In that case we have that different components are decorrelated $\langle f_{l}f_{j}\rangle=0$ if $l\neq j$. Another type of forcing is divergence free and forces components $\mathbf{e}_{x}$ and $\mathbf{e}_{z}$: it is the curl of a vector potential along $\mathbf{e}_{y}$. In that case components $\mathbf{e}_{x}$ and $\mathbf{e}_{z}$ are forced and a correlation can exists between these two components of the forcing. Numerically, we chose to add the forcing after the predictor-corrector stage of time integration, in line with the formulation of an Itô Process. We checked a posteriori that incompressibility was enforced up to the same numerical precision as in unforced simulations. Note that in studies of action minimisation in plane Poiseuille flow, both an unconstrained noise white in time and space Wan, (2013) and a noise defined as a Wiener process on the space of divergent free vector fields Wan et al., (2015); Wan and Yu, (2017), still white in time, were considered. Linear Farrell and Ioannou, 1993b and weakly non linear Ducimetière et al., (2022) studies of flows used noise with little constraint on spectrum shape and component correlations. In the large majority of cases, we will not directly force the spatial scales expected in the flow of order $\mathcal{O}(1)$ or larger: we force below a corresponding cut-off scale $L_{c}$. The value of this scale will be specified in each forcing case. What we do is that we let the flow non linearly select its scales from energy injected solely at moderately small scales (some comparison with forcing all large scales have been performed). Both types of forcing typically give a flat spectrum for the forcing at intermediate wave numbers and after a wavenumber cut-off $N_{r}$, we will use a decaying spectrum for the forcing at larger wavenumbers. We display an example of shape factor $\Gamma_{x,n_{x},n_{z}}$ in the divergence free case in figure 2 (a) and $\Gamma_{z,n_{x},n_{z}}$ in figure 2 (b), we we chose $L_{c}=1.0$. As a comparison, the spectra of $u_{x}$ and $u_{z}$ computed in the midplane $y=0$ for transitional turbulence in a MFU type domain at the largest resolution are displayed in figure 2 (c) and figure 2 (d) respectively. We can notice that these two components have their intense modes in the wavenumber range $\left|\frac{2\pi n_{x}}{L_{x}}\right|\leq 3$, $\left|\frac{2\pi n_{z}}{L_{z}}\right|\leq 6$, where the forcing is absent. The forcing starts to act in a wavenumber range where the natural Fourier modes of the velocity field are at least one order of magnitude smaller than their maximum.
@@ 2.1.2 Diagnostics of the flow
$E_{k,l}=\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{u_{l}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z\,,\\ E_{k,y-z}=\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{u_{y}^{2}+u_{z}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z\,.$ (5)
The kinetic energy in the streamwise component $E_{k,x}$ will be the most used. The energy $E_{k,x}$ roughly quantifies the amount of energy contained in the velocity streaks, while the energy $E_{k,y-z}$ roughly quantifies the amount of energy contained in the streamwise vortices, which are the two main coherent structures in transitional wall flows. These quantities have been able to highlight the processes at play in the collapse of turbulence Rolland, (2022), and are quite close to what can be measured in experiments by means of two dimensional, two components particle image velocimetry in planes parallel to the moving wall Liu et al., (2021). Note that other choices can be made to follow the amplitude of streamwise vortices. For instance isocontour of streamwise vorticity can be added to isocontour of streamwise velocity in visualisations Toh and Itano, (2003); Cherubini et al., (2011). Other choices include colour levels of streamwise vorticity Willis and Kerswell, (2009), or isocontours of the Q-criterion Rigas et al., (2021). In order to have a global measure corresponding to these visualisations, one could use the square of the streamwise component of vorticity. We stick to the kinetic energy $E_{K,y-z}$ because of the simplicity of the diagnostic and because contribution of other types of eddies is relatively small in our case. We finally used a turbulent fraction $F$ to observe the possible spatial localisation of turbulence. This indicator is computed in the following manner: we spatially average the kinetic energy contained in components $u_{y}$ and $u_{z}$ in $\tilde{N_{z}}$ streamwise bands of width $\delta z=L_{z}/\tilde{N_{z}}=2.0$, to define
@@
We now introduce a parameter alternative to $\beta$ which is possibly better suited to follow the transitions: the energy injection rate. This parameter is present in the ensemble average of the spatially average kinetic energy budget. In order to obtain this budget, we multiply equation (2) by $u_{l}$, sum over the three components and simplify using incompressibility. We perform a spatial average, to obtain the time derivative of the spatially averaged kinetic energy (Eq. (4)), and we perform ensemble average with respect to the forcing noise $\mathbf{f}$, in order to remove it from the equation. In appendix B, we detail how we use Itô’s Lemma (see (Gardiner,, 2009) § 4) to manage this non linear change of variables. This yields the average energy budget
@@
in the simulation. Note that equation (7) is similar to the Reynolds–Orr equation which has often been used for non-linear stability analysis of wall flows Drazin and Reid, (2004); Stuart, (1958). In our case we have an ensemble average of each term and the addition of the energy injection rate. In this energy budget (Eq. (7)), the extraction is performed by lift-up, in the streamwise direction, thus involving $u_{x}$ and $u_{y}$ and the wall normal gradient of the laminar baseflow equal to $1$. Dissipation is performed by viscosity. In the case of an unforced turbulent plane Couette flow, these two terms compensate one another after time average. For small energy injection rate $\epsilon$, it is generally expected that the flow remains close to laminar and that equation (2) can be treated linearly Farrell and Ioannou, 1993b or could be processed using a weakly non linear approach Ducimetière et al., (2022). However, this behaviour is conditioned on no transition happening and is thus not really represented by the actual ensemble average in equation (7). We will see in the rest of this text that even in that case, transitions can occur. This means that starting from the initial condition $\langle E_{k}\rangle(t=0)=0$ even with a small $\epsilon$, the solution of equation (7) would display a very slow but steady increase of $\langle E_{k}\rangle(t>0)$ until the flow becomes turbulent. Let us discuss how transition would be represented in this energy budget. We would expect that $\epsilon$ would act against dissipation and could maintain some energy extraction (that would not be sustainable without energy injection, regardless of the spectrum shape). It is possible that a large part of the energy transfers would not be visible in this equation as they would be erased by the spatial averages. We would also expect that the larger $\epsilon$ is, the more efficiently transition is forced, similarly to the dependence on amplitude of perturbations to boundary layers Spangler and Wells Jr, (1968); Rigas et al., (2021). The energy injection rate is controlled by two parameters whose effect on the build-up can be presented independently. Indeed, from equation (8), we deduce that $\beta$ and the spectrum shape are the two independent control parameters of this energy injection rate $\epsilon$. Thus, on the one hand, decreasing $\beta$ increases $\epsilon$, and on the other hand, increasing the sum of the $\Gamma_{l,n_{x},n_{z}}$ also increases $\epsilon$: in both cases, we force the flow more intensely. In our definitions (§ A), these shape factors $\Gamma$ will be normalised and will decrease at large wavenumbers. Thus the sum will be larger if that decrease starts at larger wave numbers and if that decrease is slow. Note that we cannot rule out that there could be effect of the precise shape of the spectrum. These complex effects would be hidden in the way energy is transferred from the forced scales to the most receptive scales. There can also be effects of a direct forcing of scales versus energy transfer. As a final remark, equation (8) reminds that in order to have a finite or even vanishing energy injection rate in cases where the flow is forced with a noise white in space (see Wan, (2013)), one would need to take the limit of $\beta\rightarrow 0$ first and $N_{x},N_{y},N_{z}\rightarrow\infty$ next (or use specific relation between those parameters). With our spectrum shape, we keep the energy injection rate finite, similarly to what is done in study of linear response Farrell and Ioannou, 1993b and non linear response Ducimetière et al., (2022) to stochastic forcing. This physical argument leads to the same mathematical condition for well posedness of the Freidlin–Wentzell Large Deviation principle in two dimensional Navier–Stokes equations as noted by Wan et al., (2015) and demonstrated by Brzeźniak et al., (2015).
@@
Note that a similar energy budget is also commonly used in the study of transition to turbulence Kawahara and Kida, (2001); Gibson et al., (2008); Kreilos and Eckhardt, (2012). This time, it is obtained for the total spatially averaged kinetic energy $E_{\rm tot}$. For this matter, we apply Itô’s Lemma for the change of variable from the Navier–Stokes equations for total flow $\mathbf{u}+y\mathbf{e}_{x}$, to the equation giving the evolution of spatially averaged total kinetic energy (see again § B). After ensemble averaging, this gives us
$\frac{dE_{\rm tot}}{dt}=\frac{1}{R}(I-D)+\epsilon\,,$
$,E_{\rm tot}=\int_{x=0}^{L_{x}}\int_{y=-1}^{1}\int_{z=0}^{L_{z}}\frac{(u_{x}+y)^{2}+u_{y}^{2}+u_{z}^{2}}{2}\,{\rm d}x{\rm d}y{\rm d}z$ (9)
$I=1+\frac{1}{2L_{x}L_{z}}\int_{x=0}^{L_{x}}\int_{z=0}^{L_{z}}\left(\left.\frac{\partial u_{x}}{\partial y}\right|_{y=+1}+\left.\frac{\partial u_{x}}{\partial y}\right|_{y=-1}\right)\,{\rm d}x{\rm d}z\,,$
@@
where $I$ is the total energy injection at the wall and $D$ is the total dissipation, written using total vorticity because of incompressibility, periodic and no slip boundary conditions. As such, both are rescaled by their value in the laminar flow where $\mathbf{u}=\mathbf{0}$ and thus $I=D=1$. This pair of variable is convenient to observe transition in wall flows. The line $I=D$ has a special importance because all invariant states lie on it (since $\frac{dE_{\rm tot}}{dt}=0$ on them), while limit cycles lie over it (since the time average over a period of $\frac{dE_{\rm tot}}{dt}$ is 00 on them). Meanwhile turbulent flow is close to $I=D=3$ Kawahara and Kida, (2001); Gibson et al., (2008). Note that streamwise vortices contribute strongly to the dissipation. However, it is difficult to disentangle their contribution from shear.